OpenStudy (anonymous):

differentiate y=cosec^-1(x^2-1)

OpenStudy (amistre64):

csc derives similar to sec; sec' = sec tan csc' = -csc cot

OpenStudy (amistre64):

but this is inverse stuff ....

OpenStudy (amistre64):

csc-1(x)=y x = csc(y) ; derive 1 = -csc(y) cot(y) y' ; solve for y' 1 ----------- = y' -csc(y) cot(y) the rest is just reforming those y parts back into xs

OpenStudy (amistre64):

since csc(y) = x thats an easy one 1 ------- = y' -x cot(y) for the cot, it might be good to build a righty: cot(y) = sqrt(x^2-1) 1 - ------------- = y' x sqrt(x^2-1)

OpenStudy (amistre64):

of course, you want to implement your argument into this :)

OpenStudy (amistre64):

i shoulda used a more generic "u" rather than an "x" prolly. lets say: u = x^2-1, then u' - ------------- = y' u sqrt(u^2-1) and so we simply rewrite the us back into (x^2-1)s

OpenStudy (anonymous):

\[-2/{(x ^{2}-1)\sqrt{(x ^{2}-2})}\] i was looking for an answer to this prob and this is what i got. want to know whether this is correct or not.

OpenStudy (amistre64):

the -2 is good the x^2-1 is good as long as: sqrt((x^2-1)^2 -1) pans out ....

OpenStudy (amistre64):

(x^2-1)^2 = x^4 -2x^2 +1 x^4 -2x^2 +1 - 1 = x^4 -2x^2 x^4 -2x^2 = x^2 (x^2-2) sqrt(x^2 (x^2-2)) = x sqrt(x^2-2)

OpenStudy (amistre64):

id say your missing an "x"

OpenStudy (anonymous):

derivative of (x^2-1) at the end gives 2x and x is cancelled.

OpenStudy (amistre64):

yay!! ... then your right ;)

OpenStudy (amistre64):

i good place to remove the human element when checking answers is: wolframalpha.com